OpenStudy (anonymous):

PLEASE EXPLAIN the difference between a removable and non-renewable discontinuity?

OpenStudy (anonymous):

non - removable?

OpenStudy (anonymous):

damn i had a decent explanation, but 15 minutes is apparently too long to wait...

OpenStudy (anonymous):

\[if \lim_{x \rightarrow a} f \left( x \right) can be made equal \to f \left( a \right) by assigning suitable value of f \left( a \right)\] ,then it is called a removable discontinuity. if it can not be made to equal to f(a) for any value of x,then it is called a non-removable discontinuity.